OpenStudy (anonymous):

***edited*** how do i do this??? 2x + 2 = 10

OpenStudy (anonymous):

put constants on one side and see what you get

OpenStudy (anonymous):

constants ??

OpenStudy (anonymous):

do u guys know?

OpenStudy (amistre64):

"what" + 2 = 10?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

then 2x has to equal 8 2 times "what" = 8 ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

good, then x has to be 4

OpenStudy (anonymous):

i get it lol thanks

OpenStudy (anonymous):

simple 2x+2=10 so you look at the 2 you are going to subtract it from itself so +2-2=0 then -2-10=8 so bring it down 2x+8 then divide so the answer is 4

OpenStudy (amistre64):

if you wanna go the algebra route: 2x + 2 = 10 ; subtract 2 from each side -2 -2 ----------- 2x = 8 ; divide off the 2 /2 /2 ----------- x = 4